### v1

# GoalRank: Group-Relative Optimization for a Large Ranking Model

Abstract
Mainstream ranking approaches typically follow a Generator–Evaluator two-stage paradigm, where a generator produces candidate lists and an evaluator selects the best one. Recent work has attempted to enhance performance by expanding the number of candidate lists, for example, through multi-generator settings. However, ranking involves selecting a recommendation list from a combinatorially large space, simply enlarging the candidate set remains ineffective, and performance gains quickly saturate. At the same time, recent advances in large recommendation models have shown that end-to-end one-stage models can achieve promising performance with the expectation of scaling laws. Motivated by this, we revisit ranking from a generator-only one-stage perspective. We theoretically prove that, for any (finite Multi-)Generator–Evaluator model, there always exists a generator-only model that achieves strictly smaller approximation error to the optimal ranking policy, while also enjoying a scaling law as its size increases. Building on this result, we derive an evidence upper bound of the one-stage optimization objective, from which we find that one can leverage a reward model trained on real user feedback to construct a reference policy in a group-relative manner. This reference policy serves as a practical surrogate of the optimal policy, enabling effective training of a large generator-only ranker. Based on these insights, we propose GoalRank, a generator-only ranking framework. Extensive offline experiments on public benchmarks and large-scale online A/B tests demonstrate that GoalRank consistently outperforms state-of-the-art methods.
1 Introduction
Recommender systems are indispensable for coping with the exponential growth of online content (Gomez-Uribe & Hunt, 2015). Industrial platforms typically adopt a multi-stage pipeline, comprising retrieval (He et al., 2020; Zhang et al., 2024) and ranking (Yu et al., 2019; Liu et al., 2023; Zhang et al., 2025). The ranking stage is particularly critical, as it determines the final sequence of items shown to users and has a major impact on both user satisfaction and platform revenue.
Formally, the ranking task can be defined as an $N{\!\rightarrow\!}L$ list-generation problem: given $N$ candidates from the preceding stage, the model outputs an ordered list of length $L$. The search space is the set of length-$L$ permutations, $\mathrm{P}(N,L)=\tfrac{N!}{(N-L)!}$, which makes exhaustive enumeration intractable for large $N$. Early approaches adopt a one-stage single generator that directly produces recommendation lists by scoring items and arranging them greedily (Zhuang et al., 2018; Ai et al., 2018; Pei et al., 2019a; Gong et al., 2022; Liu et al., 2023), as illustrated in Figure 1(a). However, this greedy strategy only models the item interdependencies in the candidate set (of size $N$) but not in the output list (of size $L$), often resulting in suboptimal rankings.
To address this limitation, subsequent studies propose a two-stage Generator–Evaluator paradigm (Shi et al., 2023; Xi et al., 2024; Lin et al., 2024; Ren et al., 2024b; Zhang et al., 2025) (Figure 1b): a generator first proposes multiple candidate lists, and an evaluator then selects the best one according to an estimated list-wise value. To mitigate the risk that generators produce only locally optimal candidates, later works introduce multi-generator settings (Figure 1c), thereby increasing both the number and diversity of candidate lists. In practice, however, simply scaling the number of candidates or generators yields diminishing returns, with performance gains quickly plateauing (Figure 1d).
Meanwhile, advances in end-to-end, one-stage large recommendation models suggest that a single sufficiently expressive model can subsume multi-stage pipelines, avoid cross-stage inconsistencies, and exhibit favorable scaling behavior (Zhai et al., 2024; Deng et al., 2025). These findings indicate that the two-stage Generator–Evaluator paradigm may not be indispensable for achieving high-quality ranking. Motivated by this, we revisit the generator-only paradigm and ask: can a larger, more powerful one-stage ranker directly produce high-quality lists without relying on an external evaluator? Formally, let $\pi^{*}$ denote the optimal ranking policy. We focus on two central questions:
  1. i
    For any (finite Multi-)Generator–Evaluator system, does there exist a single generator-only model whose policy achieves a strictly smaller approximation error with respect to $\pi^{*}$?
  2. $i$
    If such a model exists, how can it be trained effectively to realize this approximation advantage in practice?
Figure 1: Illustration of different ranking paradigms: (a) Generator-only; (b) Generator–Evaluator; (c) Multi-Generator–Evaluator; and (d) Performance trend with increasing number of generators.
To answer these questions, we analyze (in Section 3.1) the approximation error between the policy space induced by a finite set of (Multi-)Generator–Evaluator models and the optimal ranking policy $\pi^{*}$. This analysis proves the existence of a generator-only model that can achieve a strictly smaller approximation error. Moreover, we show that as the size of this generator-only ranking model increases, its approximation error with respect to $\pi^{*}$ decreases accordingly. Building on these theoretical insights, we then turn to the practical challenge of how to train such a one-stage ranking model. By deriving an evidence upper bound of the existing optimization objective, we find that one can leverage a reward model trained on real user feedback to construct a reference policy in a group-relative manner, which serves as a surrogate for $\pi^{*}$. This enables us to train a large generator-only ranking model effectively. Based on this idea, we propose a new training framework, GoalRank (Group-Relative OptimizAtion for a Large Ranker). We validate the effectiveness of GoalRank on public benchmarks as well as through large-scale online A/B tests, showing substantial improvements over state-of-the-art baselines and clear evidence of scaling laws.
The main contributions of this work can be summarized as follows:
  • Theoretical foundation. We prove that for any (finite Multi-)Generator–Evaluator family, there always exists a generator-only model that achieves a strictly smaller approximation error to the optimal ranking policy, and that this error decreases as model size increases (scaling law).
  • Optimization principle. We introduce the Group-Relative optimization principle, which provides a tractable and effective criterion for training large generator-only ranking models.
  • Model and validation. We instantiate these ideas in GoalRank, a generator-only large ranker trained under the proposed principle. Extensive offline experiments and online A/B tests demonstrate consistent improvements over strong baselines and reveal clear scaling laws with respect to model capacity.
2 Preliminary
Let $\mathcal{U}=\{u\}$ be the user set and $\mathcal{V}=\{v\}$ be the item set. For each user $u\in\mathcal{U}$, let $\mathcal{V}_{u}\subset\mathcal{V}$ denote the candidate item set, with $|\mathcal{V}_{u}|=N$. Given a list length $L<N$, the list generation space for user $u$ is given by:
$\mathcal{L}_{u}:=\bigl\{\,l=(v_{1},\dots,v_{L})\in\mathcal{V}_{u}^{L}\,\big|\,v_{i}\neq v_{j}\ (i\neq j)\,\bigr\},\quad|\mathcal{L}_{u}|=\mathrm{P}(N,L):=\frac{N!}{(N-L)!}.$
Single-Stage Ranking. (G-only) A single-stage ranker directly generates a list through a generator. Let $\mathcal{X}_{u}$ denote the context information of user $u$ (e.g., historical interactions, side features). Given a ranking model $G(\cdot)$, the single-stage ranking can be formulated as:
$l_{u}^{*}=G(\mathcal{X}_{u},\mathcal{V}_{u}).$
Two-Stage Ranking. ((M-)G-E) In contrast, a two-stage ranking model employs an evaluator $E(\cdot)$ to assess the quality of candidate lists, and then selects the best one. Formally, the task can be expressed as
$l_{u}^{*}=\arg\max_{l\in\mathcal{L}_{u,k}}E(\mathcal{X}_{u},l),\quad\text{s.t.}\quad\mathcal{L}_{u,k}=\{\,l_{i}\mid l_{i}=G_{i}(\mathcal{X}_{u},\mathcal{V}_{u}),\ i=1,\dots,k\,\}.$
The single-generator–evaluator case corresponds to $k=1$, while the multi-generator–evaluator paradigm uses $k>1$ to enlarge the candidate-list space before evaluation. Due to constraints on space, a detailed discussion of related works is provided in Appendix A.
3 Methodology
In this section, we address the research questions raised in Section 1, namely: (i) can the generator-only paradigm outperform the (Multi-)Generator–Evaluator paradigm, and (ii) if so, how can such a generator-only ranking model be effectively learned? Building on the insights gained from these analyses, we then propose a new generator-only large ranker framework, GoalRank, which leverages group-relative optimization to approximate the optimal ranking policy.
3.1 Can the generator-only paradigm perform better?
To assess the feasibility of a single-stage large ranking model, we first ask whether a sufficiently large generator-only model can match or even exceed the expressive power of the widely used two-stage (Multi-)Generator-Evaluator pipeline. Formally, suppose there exists an optimal ranking policy $\pi^{*}$. We compare the best attainable approximation error of (i) a $k$-mixture of small generators combined with an evaluator, and (ii) a single larger generator. To make this comparison precise, we begin by defining a capacity-restricted generator class.
Definition 1 ($(\alpha,\beta)$-bounded generator class).
Given maximum generator width $\alpha$ and depth $\beta$, the $(\alpha,\beta)$-bounded generator class is defined as
$\mathcal{G}_{m}(\alpha,\beta)\;:=\;\bigl\{\,g_{m}\;\bigm|\;W(g_{m})\leq\alpha,\;\;D(g_{m})\leq\beta\,\bigr\},$
where $g_{m}$ denotes a generator, and $W(\cdot)$ and $D(\cdot)$ measure width- and depth-type complexities, respectively.
Then, the evaluator can be regarded as operating over a low-dimensional probability simplex, which determines how multiple small generators jointly influence the final ranking policy.
Definition 2 ($k$-mixture $(\alpha,\beta)$-bounded policy space).
The policy space induced by $\mathcal{G}_{m}(\alpha,\beta)$ is
$\mathcal{F}_{m}(\alpha,\beta)\;:=\;\bigl\{\,\mathrm{softmax}\!\circ g_{m}\;\bigm|\;g_{m}\in\mathcal{G}_{m}(\alpha,\beta)\,\bigr\},$
which contains all policies realizable by a single generator in $\mathcal{G}_{m}(\alpha,\beta)$ with a softmax output layer. Given $k$ generators in $\mathcal{G}_{m}(\alpha,\beta)$ and an evaluator, the corresponding $k$-mixture $(\alpha,\beta)$-bounded policy space is
$\mathcal{C}_{m}^{k}(\alpha,\beta)\;:=\;\Bigl\{\,\textstyle\sum_{i=1}^{k}\omega_{i}\,\pi_{i}\;\Bigm|\;\bm{\omega}\in\Delta^{k-1},\;\pi_{i}\in\mathcal{F}_{m}(\alpha,\beta)\Bigr\},$
where $\Delta^{k-1}$ is the (k1)(k-1)-dimensional probability simplex and $\bm{\omega}=(\omega_{1},\ldots,\omega_{k})$ satisfies $\sum_{i=1}^{k}\omega_{i}=1$ and $\omega_{i}\geq 0$.
In Definition 2 we adopt soft mixture weights $\bm{\omega}$. In practice, the evaluator often implements (or approximates) a hard selection (one-hot $\bm{\omega}$). Thus, $\mathcal{C}_{m}^{k}(\alpha,\beta)$ strictly contains the policy class realized by hard selection, which can both simplifies subsequent derivations and strengthens Theorem 1. Then, to evaluate how well a policy space approximates the optimal ranking policy $\pi^{*}$, we use the following notion.
Definition 3 (Approximation distance (KL error)).
Let $\pi^{*}$ be a target policy and $\mathcal{F}$ be a policy space. The approximation distance from $\mathcal{F}$ to $\pi^{*}$ is
$\mathcal{E}(\mathcal{F})\;:=\;\inf_{\pi\in\mathcal{F}}\mathrm{KL}\!\bigl(\pi^{*}\|\pi\bigr),\qquad\mathrm{KL}\!\bigl(\pi^{*}\|\pi\bigr)\;=\;\sum_{l\in\mathcal{L}}\pi^{*}(l)\,\log\!\frac{\pi^{*}(l)}{\pi(l)}.$
where $\mathcal{L}$ denotes the finite space of candidate lists considered by the ranker.
With these definitions in place, we can now state our main result.
Theorem 1.
Given $\alpha,\beta>0$ and any $k\in\mathbb{N}_{>0}$. For the $k$-mixture policy space $\mathcal{C}_{k}^{m}(\alpha,\beta)$ in Definition 2, there exists a class of larger generators
$\mathcal{G}_{M}(\alpha,\beta,n)\;:=\;\bigl\{\,g_{M}\;\bigm|\;W(g_{M})\,\geq\,k\alpha+n,\;\;D(g_{M})\,\geq\,\beta\,\bigr\},\qquad n\in\mathbb{N}_{>0},$
with associated policy space
$\mathcal{F}_{M}(\alpha,\beta,n)\;:=\;\bigl\{\,\mathrm{softmax}\!\circ g_{M}\;\bigm|\;g_{M}\in\mathcal{G}_{M}(\alpha,\beta,n)\,\bigr\},$
such that
$\mathcal{E}\bigl(\mathcal{F}_{M}(\alpha,\beta,n)\bigr)\;<\;\mathcal{E}\bigl(\mathcal{C}_{m}^{k}(\alpha,\beta)\bigr),\qquad\lim_{n\to\infty}\,\mathcal{E}\bigl(\mathcal{F}_{M}(\alpha,\beta,n)\bigr)\;=\;0.$
Proofs and technical details are deferred to Appendix B. Theorem 1 shows that for any two-stage ranking mixing $k$ small generators, there exists a sufficiently large one-stage generator-only ranking model whose induced policy space achieves a strictly smaller approximation error to $\pi^{*}$. Moreover, as the size of this generator increases (i.e., as $n$ grows), the approximation error can be driven arbitrarily close to zero. We remark that Theorem 1 is stated in terms of width scaling; the same conclusion holds under depth scaling, with proofs provided in Appendix B.
3.2 how can generator-only ranking model be effectively learned?
According to Theorem 1, our goal is to train a larger generator-only ranking model that can achieve a closer approximation to the optimal ranking policy $\pi^{*}$. Suppose we have access to an ideal reward model $r^{*}(l)$ that provides an unbiased estimate of the user feedback for any candidate list $l\in\mathcal{L}_{u}$ of user uu [^1] Here, the reward value refers to the user’s actual feedback to a list, e.g., watch time or interaction behaviors.. We define the entropy-regularized oracle policy as
$\pi^{*}:=\arg\max_{\pi}\Bigl\{\mathbb{E}_{l\sim\pi}\!\left[r^{*}(l)\right]+\tau\mathcal{H}(\pi)\Bigr\},$ (1)
where $\mathcal{H}(\pi)$ denotes the entropy of the policy, introduced as a regularization term to avoid greedy instability and to encourage exploration, and $\tau>0$ controls the strength of entropy regularization. Optimizing Equation 1 yields the Boltzmann distribution
$\pi^{*}(l)=\frac{\exp(r^{*}(l)/\tau)}{Z},\qquad Z=\sum_{l^{\prime}}\exp\!\left(r^{*}(l^{\prime})/\tau\right).$ (2)
Moreover, the objective in Equation 1 can be equivalently rewritten as
$\mathbb{E}_{l\sim\pi}\!\left[r^{*}(l)\right]+\tau\mathcal{H}(\pi)$ $=\tau\sum_{l}\pi(l)\Bigl(\log\tfrac{\exp(r^{*}(l)/\tau)}{Z}-\log\pi(l)+\log Z\Bigr).$
Thus,
$\tau\log Z=\sup_{\pi}\Bigl\{\mathbb{E}_{l\sim\pi}\!\left[r^{*}(l)\right]+\tau\mathcal{H}(\pi)\Bigr\},$
and the supremum is attained if and only if $\mathrm{KL}(\pi\|\pi^{*})=0$. In this case, the objective achieves its maximum. Therefore, optimizing $\pi$ is equivalent to minimizing the KL divergence to $\pi^{*}$.
In practice, however, the ideal reward model $r^{*}(l)$ is inaccessible. We therefore consider a potentially biased reward model $\hat{r}(l)=r^{*}(l)+b(l)$, where $b(l)$ denotes the bias. Intuitively, the smaller the bias, the more reliable the reward model. When considering a list group $\mathcal{B}$, if the reward gaps among lists are sufficiently large, the contribution of $r^{*}(l)$ dominates the bias $b(l)$, such that the (partial) order over $\mathcal{B}$ is approximately preserved. Formally, given a threshold $\sigma^{*}>0$, if
$\max_{l_{i},l_{j}\in\mathcal{B}}\big|\hat{r}(l_{i})-\hat{r}(l_{j})\big|>\sigma^{*},$ (3)
we can exploit this approximate order-invariance to construct a reference policy in a group-relative manner:
$\pi^{\mathrm{ref}}(l\mid\mathcal{B})=\frac{\exp\!\bigl((\hat{r}(l)-\bar{r}_{\mathcal{B}})/\sigma_{\mathcal{B}}\bigr)}{\sum_{l^{\prime}}\exp\!\bigl((\hat{r}(l^{\prime})-\bar{r}_{\mathcal{B}})/\sigma_{\mathcal{B}}\bigr)},$ (4)
where $\bar{r}_{\mathcal{B}}$ and $\sigma_{\mathcal{B}}$ are the mean and standard deviation of $\hat{r}$ over $\mathcal{B}$.
We then train a parametric policy $\pi_{\theta}$ to align with $\pi^{\mathrm{ref}}$ by minimizing the cross-entropy (equivalently, the KL divergence up to a constant): Finally, the training objective can be expressed as
$\mathcal{L}(\pi_{\theta})=-\mathbb{E}_{\mathcal{B}\sim\mathcal{D}}\Biggl[\sum_{l\in\mathcal{B}}\pi^{\mathrm{ref}}(l\mid\mathcal{B})\log\pi_{\theta}(l)\Biggr].$ (5)
This objective provides a tractable surrogate for minimizing $\mathrm{KL}(\pi_{\theta}\|\pi^{*})$ using only $\hat{r}$ and group-relative normalization.
3.3 GoalRank
Based on the above insights, we propose a practical training framework for large ranking models in real-world recommendation scenarios.
Reward modeling.
Following Zhang et al. (2025), we first train a reward model $\hat{r}$ using real user feedback data, which can estimate the expected feedback for a given recommendation list (details are provided in Appendix C).
Figure 2: Training pipeline of group-relative optimization for a large ranker, GoalRank.
Generator and policy.
As illustrated in Figure 2, given a generator $g_{\theta}$, the corresponding ranking policy is defined as $\pi_{\theta}:=\mathrm{softmax}\!\circ g_{\theta}$. Conditioned on user context $\mathcal{X}_{u}$ and the candidate item set $\mathcal{V}_{u}$ provided by the preceding stage, the generator produces a recommendation list as
$l_{u}^{\theta}=\arg\max_{l}\pi_{\theta}(l\mid\mathcal{X}_{u},\mathcal{V}_{u}).$ (6)
Note that this framework is model-agnostic: the generator can be instantiated by any sequence generation model.
Group construction.
As discussed in the previous section, constructing effective groups requires sufficiently large reward gaps among lists within each group, which is difficult to achieve when sampling multiple lists from a single generator. To address this, we introduce an auxiliary set of ranking policies $\mathcal{M}$ (including heuristic methods and lightweight neural models with implementation details provided in Appendix D). For each user $u$, we then construct a group of recommendation lists as
$\mathcal{B}_{u}=\{l_{u}^{\theta}\}\cup\{l_{u}^{i}\mid l_{u}^{i}=\arg\max_{l}\pi_{i}(l\mid\mathcal{X}_{u},\mathcal{V}_{u}),\;\pi_{i}\in\mathcal{M}\}.$
As an additional option, all lists in $\mathcal{B}_{u}$ can be ranked by their rewards, and a uniformly sampled subset can then be selected, which further enforces larger reward gaps within the group and strengthens the validity of the condition in Equation 3.
Training.
Given $\mathcal{B}_{u}$, we compute the reference policy $\pi^{\mathrm{ref}}(\cdot\mid\mathcal{B}_{u})$ via Equation 4 and optimize $g_{\theta}$ by minimizing the loss in Equation 5, instantiated with user-specific groups $\{\mathcal{B}_{u}\}_{u\in\mathcal{U}}$. This realizes the group-relative principle and provides a practical path to align $\pi_{\theta}$ with the oracle policy structure using accessible signals.
4 Experiment
In this section, we present both offline and online experiments to evaluate the effectiveness of GoalRank, which are designed to address the following research questions:
  • RQ1: How does GoalRank perform on $N{\!\rightarrow\!}L$ ranking tasks compared with state-of-the-art baselines, and does it exhibit scaling behavior as model or data size increases?
  • RQ2: How do (i) the size of recommendation list group $\mathcal{B}$ and (ii) the reward model’s prediction bias affect GoalRank performance?
  • RQ3 (online): How does GoalRank perform in real-world industrial recommendation scenarios?
4.1 Offline Experiments
4.1.1 Datasets and Offline Experiments Setting
We conduct offline experiments on two public datasets, ML-1M (Harper, 2015) and Amazon-Book (McAuley et al., 2015), as well as two datasets of different scales collected from our industrial short-video platform, denoted as Industry and Industry-0.1B. The statistics of the four preprocessed datasets are summarized in Appendix E.1.
For dataset construction, we first perform an 80/20 temporal split. For each user’s interaction history (sorted chronologically), the task is framed as an $N{\!\rightarrow\!}L$ list-generation ranking problem with $N=50$ and $L=6$. Specifically, we use a pre-trained Matrix Factorization (MF) model (Koren et al., 2009) as the retriever to select the top-50 candidate items for each user. The last six interactions in each user’s historical sequence are treated as ground truth, representing the target list after ranking. For industry datasets, we define a long view (watching a video for more than 85% of its duration) as a positive signal, indicating meaningful user–item engagement.
Following common practices, we report Hit Ratio@L (H@L), NDCG@L (N@L), MAP@L (M@L), F1@L, and AUC with $L=6$. All reported results are averaged over five independent runs.
4.1.2 Baselines
We compare GoalRank against representative state-of-the-art methods from three major categories of ranking models:
  • Generator-only methods: These approaches rely on a single generator to produce item scores and directly generate the ranking list. Simple item-wise scoring models such as DNN (Covington et al., 2016) estimate user feedback independently for each user–item pair. More advanced methods, including DLCM (Ai et al., 2018), PRS (Feng et al., 2021a), PRM (Pei et al., 2019a), and MIR (Xi et al., 2022), explicitly capture mutual dependencies among candidate items.
  • Generator–Evaluator methods: These methods (e.g., EGRerank (Huzhang et al., 2021), PIER (Shi et al., 2023), and NAR4Rec (Ren et al., 2024b)) first generate multiple reranked candidate lists and then leverage an evaluator to select the most effective one for the user. Following Yang et al. (2025), for PIER, we first apply a pointwise ranking model to select the top-6 items, enumerate all possible permutations, and then use the evaluator to identify the optimal ranking.
  • Multi-Generator–Evaluator methods: These approaches ensemble multiple generators to expand the candidate-list space and enhance ranking performance (Yang et al., 2025). In our experiments, we evaluate this strategy under different ensemble sizes, with the number of generators set to 3, 20, and 100.
To ensure fairness, all baselines are tuned within their respective parameter spaces. Unless otherwise specified (e.g., in scaling law experiments), the hidden embedding dimension of all models is fixed at 128, and model depths are kept consistent. Further details are provided in Appendix E.2.
4.1.3 Main Results (RQ1)
Table 1: Overall performance of different ranking methods. The highest scores are in bold, and the runner-ups are with underlines. All improvements are statistically significant with student t-test p<0.05p<0.05.“Improv.” denotes the improvements over the best baselines.
Methods ML-1M Industry Book
H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC
G-only DNN 56.86 70.30 59.28 62.16 86.87 37.32 54.56 42.38 43.72 74.73 60.28 69.61 58.58 62.45 83.02
DLCM 62.31 73.87 63.82 67.96 89.35 39.69 60.67 48.90 46.61 75.80 66.80 75.88 65.39 69.28 91.93
PRS 59.35 73.10 62.51 64.72 88.84 44.75 64.39 51.88 52.59 89.93 66.15 75.70 64.84 68.64 92.01
PRM 60.09 72.85 62.21 65.51 88.20 39.92 55.93 42.97 46.18 85.15 67.86 76.88 66.44 70.42 92.00
MIR 62.22 74.33 64.47 67.97 87.76 37.01 55.79 43.16 44.50 79.95 66.08 71.48 56.62 68.62 91.82
G-E EGRank 62.76 74.75 64.97 68.46 88.72 40.09 59.01 47.52 47.06 77.44 70.73 80.75 72.40 73.33 89.40
PIER 62.74 75.99 65.98 68.74 90.43 45.35 65.11 52.55 53.35 90.93 71.14 80.22 71.62 73.74 92.26
NAR4Rec 62.81 75.01 65.42 68.31 88.30 44.31 63.83 51.45 52.08 89.94 70.08 79.46 70.69 72.66 92.44
MG-E G-3 55.51 67.39 55.52 55.51 60.73 49.42 68.29 56.23 55.50 83.44 68.76 76.36 65.82 71.33 85.44
G-20 58.66 69.86 58.60 64.18 81.76 52.66 70.70 59.02 61.81 76.46 72.99 78.68 68.66 75.72 77.07
G-100 60.64 70.97 59.93 66.29 76.48 55.77 72.35 60.95 64.27 75.30 77.21 82.15 73.78 80.09 77.36
[…]
Improv. +17.12% +9.79% +15.43% +16.60% +7.98% +25.39% +20.15% +29.63% +28.04% +7.85% +4.07% +3.32% +5.60% +4.18% +2.19%
Ranking Performance. Table 1 reports the overall results across three datasets. We highlight three key observations:
  • GoalRank consistently achieves the best performance. Across all datasets and metrics (H@6, N@6, M@6, F1@6, and AUC), GoalRank outperforms all baselines. On ML-1M, it improves H@6 and M@6 by +17.12% and +15.43%; on the large Industry dataset, the gains are even more pronounced, reaching +25.39% in H@6 and +29.63% in M@6. These results confirm that a single large generator can better capture global ranking signals than multi-stage alternatives.
  • Two-stage G-E methods outperform early G-only approaches. Models such as PIER and NAR4Rec surpass DNN, DLCM, and PRM by explicitly modeling list-wise utility, validating the effectiveness of evaluators over greedy generation.
  • MG-E further improves but quickly saturates. As generator count increases from 3 to 100, H@6 rises from 49.42 to 55.77 on Industry. However, the gains diminish rapidly, and even the strongest MG-E models remain far below GoalRank, underscoring the inefficiency of enlarging the candidate set alone.
In short, while G-E and MG-E offer incremental gains, GoalRank’s generator-only design provides a more scalable and fundamentally stronger solution.
Figure 3: Scaling performance of GoalRank and baselines on the Industry-0.1B dataset across model sizes from 1M to 0.1B parameters.
Scaling Performance. In Section 3.1, we showed theoretically that GoalRank admits scaling laws. Here, we empirically validate this by varying hidden dimensions, layer depth, and attention heads, adjusting model size from 1M to 0.1B. We compare GoalRank with three representative baselines: DNN, PIER, and MG-E. For fairness, DNN and PIER are scaled in the same manner as GoalRank, while the size of MG-E is increased by enlarging the number of generators.
Figure 3 presents the scaling performance on the Industry-0.1B dataset with respect to Hit Ratio@6 (H@6), NDCG@6 (N@6), MAP@6 (M@6), and F1@6. We exclude AUC since GoalRank already achieves values above 0.98 even at small model sizes, though we observe further improvements as the model size increases, approaching 1.0. The model size ranges from 1M to 0.1B parameters. [^2] For very small models, training on the full dataset leads to unstable convergence. To ensure fair comparison, we proportionally sample the dataset for all models (including GoalRank) at the same parameter scale. We summarize the key findings as follows:
  • GoalRank demonstrates strong scaling. Metrics improve steadily from 1M to 0.1B, with the sharpest gains between 10M and 0.1B, confirming clear scaling laws.
  • DNN and PIER show weak scaling. Larger sizes yield only modest improvements, reflecting the inherent limitation of pointwise scoring in approximating the optimal policy.
  • MG-E saturates. Adding generators helps initially but plateaus quickly, indicating diminishing returns compared with GoalRank’s single-model scaling.
4.1.4 Ablation Study (RQ2)
In this section, we present ablation studies to examine two key factors: (i) the impact of the group size $|\mathcal{B}|$ on constructing the reference policy, and (ii) the robustness of GoalRank under varying levels of bias in the reward model. We report results on the Industry dataset, which is representative of the overall trends observed across all datasets.
Table 2: Performance with varying group sizes $|\mathcal{B}|$.
$|\mathcal{B}|$ 3 5 8 10 20 50 80 100
H@6 62.88 64.52 69.95 69.94 69.93 67.34 63.29 63.50
N@6 81.76 83.42 86.82 87.17 86.93 85.33 82.76 82.94
M@6 72.50 74.44 78.86 79.34 79.01 76.96 73.69 73.92
F1@6 74.00 75.94 82.25 82.28 82.29 79.26 74.47 74.76
AUC 97.50 97.79 98.05 98.15 98.07 98.06 97.75 97.77
Table 3: Performance with varying bias level $\lambda$.
$\lambda$ H@6 N@6 M@6 F1@6 AUC
0.0 69.93 86.93 79.01 82.29 98.07
0.2 65.32 83.21 74.33 76.89 97.90
0.5 63.77 82.75 73.79 75.05 97.73
Influence of the Size of $\mathcal{B}$. Table 3 reports the performance of GoalRank under different group sizes $|\mathcal{B}|$. We observe that very small groups (3–5) fail to provide enough samples for constructing a reliable reference policy, while overly large groups (50–100) weaken the reward gaps mentioned in Equation 3 and thus amplify the bias of the reward model. The best performance is achieved with moderate group sizes (8–20), which strike a balance between sample sufficiency and bias mitigation. Importantly, GoalRank consistently outperforms the best baseline even when $|\mathcal{B}|$ is set suboptimally.
Influence of Prediction Bias of Reward Models. As discussed in Section 3.2, the reward model $\hat{r}$ used to construct the reference policy is inevitably biased. To examine the effect of this bias, we introduce controlled noise by defining
$\hat{r}_{\text{bias}=\lambda}(l):=(1-\lambda)\hat{r}(l)+\lambda\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$
and evaluate GoalRank under different bias levels $\lambda\in\{0.0,0.2,0.5\}$. Results in Table 3 show that performance degrades only slightly as $\lambda$ increases, indicating that GoalRank is robust to reward model bias. Remarkably, even with $\lambda=0.5$, GoalRank still outperforms state-of-the-art baselines.
4.2 Live Experiments
4.2.1 Experimental Setup
We evaluate GoalRank through a large-scale online A/B test on a real-world short-video recommendation platform. The platform serves over half a billion daily active users, with an item pool of tens of millions of videos. The system follows a two-stage workflow: (i) retrieval, which selects $N$ candidate items from millions, and (ii) ranking, which generates a final recommendation list of length $L$. GoalRank is deployed in the ranking stage, where $N=120$ candidates are provided and $L=6$ items are exposed to each user. The overall workflow is illustrated in Figure 4 (Appendix).
4.2.2 Evaluation Protocol
We randomly partition online traffic into eight buckets, each covering about one-eighth of total users (tens of millions per bucket). We compare three settings: the production MG-E baseline(which consists of tens of generator models and hundreds of candidate lists), a hybrid setting where GoalRank serves 30% of the traffic alongside MG-E, and a pure GoalRank deployment. Each A/B test runs for at least 14 days to ensure statistical reliability. We track standard business metrics, including App Stay Time (a key overall engagement indicator), Watch Time (average continuous viewing length), Effective Views (total view count), and other behavior-specific rates on recommended items.
4.2.3 Results Analysis
Table 4: Online performances improvement of GoalRank. All results are statistically significant.
Method APP Stay Time Watch Time Effective View Like Comment
[…]
We report the results of the two-week online A/B tests in Table 4. GoalRank consistently outperforms the production MG-E framework across all business-critical metrics. Even in the hybrid setting (GoalRank + MG-E), we observe significant gains, while a full deployment of GoalRank yields the largest improvements. These results demonstrate that GoalRank can not only complement but also fully replace the existing MG-E framework, providing superior ranking quality without incurring trade-offs. GoalRank + MG-E has been deployed to serve the full user traffic in production.
5 Conclusion
In this work, we revisit the design of ranking models in recommender systems and challenged the prevailing (Multi-)Generator–Evaluator paradigm. We theoretically proved that, for any (finite Multi-)Generator–Evaluator model, there always exists a generator-only ranker that achieves strictly smaller approximation error to the optimal ranking policy, and that this error decreases further as model size grows. Building on this result, we derived an evidence upper bound of the one-stage objective and introduced the group-relative optimization principle, which leverages a reward model trained on real user feedback to construct a reference policy and provides a practical training objective for generator-only rankers. We instantiated these insights in GoalRank, a large generator-only ranker optimized under the proposed principle. Extensive offline and online experiments demonstrated that GoalRank consistently outperforms SOTA methods and exhibits clear scaling behavior.
Limitation and Future Work. In real-world applications, ranking often needs to accommodate diverse and frequently changing business objectives. Compared with (Multi-)Generator–Evaluator models, a generator-only framework like GoalRank is less flexible in adapting to such shifts. A promising direction is to incorporate business-specific contextual signals into GoalRank, thereby enhancing its adaptability and generalization across objectives. Moreover, recent progress in large recommendation models has demonstrated remarkable success in the retrieval stage. However, most of these efforts overlook list-wise modeling, limiting their ability to capture the benefits unique to ranking. Future work may therefore explore how large retrieval and large ranking models can be jointly optimized to build more powerful end-to-end recommender systems.
For the theoretical results, detailed derivations are provided in Appendix B. For the empirical studies, we will release the implementation and training code at https://github.com/Kaike-Zhang/GoalRank to ensure reproducibility.
Appendix A Related Work
The ranking task in recommender systems can be viewed as an $N{\!\rightarrow\!}L$ list-generation problem: given $N$ candidate items, the model outputs an ordered list of length $L$. Since the search space $\mathrm{P}(N,L)=\tfrac{N!}{(N-L)!}$ is combinatorially large, a series of paradigms have been developed to approximate this process.
One-stage generator-only methods. Early approaches adopt a single generator that directly produces ranking lists by scoring items and arranging them greedily (Zhuang et al., 2018; Ai et al., 2018; Pei et al., 2019a; Gong et al., 2022; Liu et al., 2023; Feng et al., 2021b; Xi et al., 2022; Pei et al., 2019b). Representative models include the GRU-based DLCM (Ai et al., 2018) and Transformer-based PRM (Pei et al., 2019a), which refine scores using listwise features. While simple and efficient, these methods largely ignore interdependencies among items or suffer from inconsistencies with the initial candidate conditioning, leading to suboptimal results.
Two-stage Generator–Evaluator (G–E) methods. To address these issues, subsequent works propose a two-stage Generator–Evaluator paradigm (Chen et al., 2022; Shi et al., 2023; Xi et al., 2024; Lin et al., 2024; Ren et al., 2024b; Wang et al., 2025b). In this framework, the generator proposes multiple candidate lists, and the evaluator selects the best one by estimating listwise utility. Compared with heuristic permutation strategies, generative approaches are more effective at exploring the exponentially large space while incorporating user preferences, and thus have become the mainstream solution.
Multi-Generator–Evaluator extensions. More recent studies extend the G–E paradigm by employing multiple generators to enlarge the candidate-list space and enhance ranking performance (Yang et al., 2025). However, the improvement is substantial only when the number of generators is small, and quickly diminishes as more generators are added. This indicates that merely scaling the number of generators is insufficient, motivating the search for fundamentally stronger rankers.
Other directions. Beyond these paradigms, recent work has explored the use of large language models (Ren et al., 2024a; Gao et al., 2024; Wu et al., 2024; Gao et al., 2025; Ren et al., 2025; Liu et al., 2025) and reinforcement learning (Feng et al., 2021c; Wang et al., 2024; 2025c; Wei et al., 2020) to improve ranking. These methods incorporate textual side information or treat ranking as a critic-based paradigm, decomposing listwise value estimation to better align recommendations with user preferences.
Appendix B Proof of Theorem 1
The detailed proof will be made publicly available after the paper is accepted for publication.
Appendix C Reward Model Training
Following Zhang et al. (2025), we train a reward model to approximate user feedback on exposed recommendation lists. Let $u\in\mathcal{U}$ denote a user with context information $\mathcal{X}_{u}$ (e.g., historical interactions or side features). Suppose $l_{u}=(v_{1},\dots,v_{|l|})$ is the exposure list presented to user $u$, where $v_{i}\in\mathcal{V}_{u}\subseteq\mathcal{V}$ and $|l|$ is the list length. The corresponding real user feedback, such as watch time, clicks, or other engagement signals, is denoted by $r_{l_{u}}\in\mathbb{R}$.
The reward model is defined as a function
$\hat{r}:\mathcal{X}\times\mathcal{V}^{|l|}\;\rightarrow\;\mathbb{R},$
which, given the user context $\mathcal{X}_{u}$ and a candidate list $l_{u}$, predicts the expected feedback $\hat{r}(l_{u}\mid\mathcal{X}_{u})$.
To train the reward model, we minimize the mean squared error between the predicted feedback $\hat{r}(l_{u}\mid\mathcal{X}_{u})$ and the observed feedback $r_{l_{u}}$ across all users:
$\mathcal{L}(\hat{r})=\mathbb{E}_{u\in\mathcal{U}}\Bigl[\bigl(\hat{r}(l_{u}\mid\mathcal{X}_{u})-r_{l_{u}}\bigr)^{2}\Bigr].$
Appendix D Methods for Group Construction
We consider multiple strategies for constructing groups of candidate lists, beyond the standard autoregressive sampling approach. The main methods are summarized as follows:
  • Autoregressive list generation (Jayaram & Thickstun, 2021): The conventional approach samples a single list by generating the entire trajectory in an autoregressive generator. While effective in capturing dependencies, this method is relatively slow and produces only one list per sampling trajectory.
  • Parallel tree-structured generation (Jayaram & Thickstun, 2021; Wang et al., 2025a): To improve efficiency and diversity, we allow the autoregressive generator to branch out in the first $K$ steps, forming a tree of partial sequences. The remaining LKL-K steps are then completed deterministically, enabling parallel exploration of multiple candidate lists.
  • Softmax-based stochastic sampling (Holtzman et al., 2019; Efraimidis & Spirakis, 2006): After the first step of scoring by generator, items are sampled probabilistically according to the softmax distribution of their scores, instead of deterministically selecting the top item, which encourages more diverse list generation.
  • Markov process approximation (Metropolis et al., 1953): The list generation process is modeled as a Markov chain, where each step only conditions on the immediately preceding item rather than the full history. We exhaustively explore and score all possible two-item pairs in the first two steps and the remaining items in the list are sampled sequentially in a chain-like manner.
  • Random selection: As a simple baseline, we randomly sample six items to form a candidate list without using model guidance.
  • Heuristic substitution (Wang et al., 2025b): Starting from already sampled lists, we heuristically replace up to two items to create new candidate lists while maintaining partial consistency with existing ones.
  • Diversity-oriented generation (Yang et al., 2025) : we generate lists that are explicitly encouraged to differ significantly from previously sampled lists, thereby enhancing the diversity of the candidate set.
Appendix E Details of Offline Experiments Settings
E.1 Details of Dataset
The ML-1M dataset is a widely used public benchmark in recommender systems, containing approximately 1 million ratings provided by over 6,000 users on more than 3,900 movies. The Amazon Books dataset is a large-scale collection of product reviews focused on books available on Amazon, consisting of about 2 million reviews from over 35,000 users spanning more than 39,000 books. The Industry dataset is collected from a real-world short-video platform that serves over half a billion daily active users and hosts an item pool of tens of millions of videos. We construct two versions of this dataset: a smaller one with over 3 million interactions from 89,310 users on 10,395 videos, and a larger one with over 0.1 billion interactions from 1 million users on 0.3 million videos.
For preprocessing, all interactions are organized in chronological order, and users/items with fewer than 20 interactions are filtered out following the standard 20-core protocol. We employ a Matrix Factorization (MF) model as the retriever simulator to generate candidate items for the ranking stage. The data is split into training and testing sets with a ratio of 8:2. For the ranking stage, interactions are sorted chronologically, and the last six interactions are used as the item list exposed to users after reranking. Table LABEL:dataset_stage1 reports the statistics of the processed datasets, including the number of users, items, interactions, and revealed lists.
Table 5: Dataset Statistics
Dataset $\left|\mathcal{U}\right|$ $\left|\mathcal{I}\right|$ # Interaction # List
ML-1M 6,020 3,043 995,154 161,646
Amazon-Book 35,732 38,121 1,960,674 311,386
Industry 89,310 10,395 3,270,132 513,010
Industry-0.1B 1,146,032 312,573 100,269,812 16,099,612
E.2 Details of Baselines
Figure 4: Online workflows.
We detail the compared baselines of our main experiments in the following, covering three major categories: Generator-only methods, Generator–Evaluator methods, and Multi-Generator–Evaluator methods:
Generator-only methods: These approaches directly predict the item-wise scores for candidates and rank them accordingly, without explicit evaluation of whole lists.
  • DNN (Covington et al., 2016) learn the user feedback for each user-item interaction.
  • DLCM (Ai et al., 2018) refines initial rankings by leveraging local context from top-retrieved documents , using a recurrent neural network to capture document interactions and an attention-based loss function to capture item interactions.
  • PRS (Feng et al., 2021a) also known as SetRank, which is a neural learning-to-rank model employs permutation-invariant neural ranking with multi-head self-attention to model cross-item dependencies, achieving robust performance across variable-length input sets.
  • PRM (Pei et al., 2019a) addresses personalized re-ranking by integrating user-specific preferences into the re-ranking process, thus enhancing both personalization and relevance.
  • MIR (Xi et al., 2022) captures complex hierarchical interactions between user actions and candidate list features to improve the accuracy of list-wise recommendation.
Generator-Evaluator Methods: These methods adopt a two-stage paradigm where the generator produces candidate lists and the evaluator selects the most promising one.
  • EGRerank (Huzhang et al., 2021) proposes an evaluator–generator framework for e-commerce ranking. The evaluator estimates list utility given context, while the generator leverages reinforcement learning to maximize evaluator scores, with an additional discriminator ensuring evaluator generalization.
  • PIER (Shi et al., 2023) follows a two-stage architecture consisting of a Fine-grained Permutation Selection Module (FPSM) and an Omnidirectional Context-aware Prediction Module (OCPM). The FPSM leverages SimHash to identify the top-$K$ candidate permutations based on user interests, while the OCPM evaluates these permutations through an omnidirectional attention mechanism.
  • NAR4Rec (Ren et al., 2024b) introduces a non-autoregressive generative re-ranking model that alleviates data sparsity and candidate variability via contrastive decoding and unlikelihood training, while also considering its integration into broader generator–evaluator frameworks.
Multi Generator-Evaluator: These methods extend the generator–evaluator paradigm by ensembling multiple generators to enlarge the candidate list space and improve final ranking quality. For example, MG-E (Yang et al., 2025) aggregates outputs from multiple generators, each specializing in different candidate distributions, before applying evaluation for list selection.